# On the embeddings of selfadjoint operator spaces

### v1

@@ 1.2. The embedding problem
This example exhibits that embeddings should reflect both the extrinsic and intrinsic aspects of positivity inherited from the C*-cover, e.g., extensions of maps, spectrum, positive decomposition and distance from cones.
@@
Theorem A. ([19], [9], Theorem 3.3, Theorem 3.8) Let $\vartheta\colon E\to F$ be a completely isometric complete order embedding between selfadjoint operator spaces. Then the following are equivalent:
@@
Equivalence [(i) $\Leftrightarrow$ (ii)] is shown in [9, Corollary 2.3]. Equivalence [(i) $\Leftrightarrow$ (iii)] is shown in a broader context in [19, Lemma 4.2 (a)]. Equivalence [(iii) $\Leftrightarrow$ (iv)] follows by a BW-compactness argument; we provide a proof that passes through Russell’s gauge maximality of item (v) and van Dobben de Bruyn’s [10] description of the positive cones of the unitisation. A key step in this approach is the distance formula
@@ 1.3. The automatic embedding property
Theorem B. (Theorem 4.5, Proposition 5.4) Let $E\subseteq\mathrm{C}^{*}(E)$ be a selfadjoint operator space. Then the following hold:
@@
Approximately positively generated spaces $E$ are in abundance and have a good position within a $\mathrm{C}^{*}(E)$. In Proposition 4.3 we combine the notion of rigidity at zero of Salomon [24] with the works of Humeniuk–Kennedy–Manor [12] and of the third author [17] to show that $E$ is approximately positively generated if and only if $M_{n}(E)$ is approximately positively generated for all $n\in\mathbb{N}$. This is an extension of [12, Proposition 8.4] to the approximation setting.
@@
Ng proves in [20, Theorem 3.9] that a selfadjoint operator space $E$ is completely $\kappa$-generated for some $\kappa>0$ if and only if it is completely approximately $\kappa^{\prime}$-generated for some $\kappa^{\prime}>0$ if and only if it is dualisable, i.e., $E^{d}$ is a selfadjoint operator space under the canonical dual matrix cone and with a matrix norm equivalent to the cb-norm. Humeniuk–Kennedy–Manor show in [12, Theorem 5.8] that dualisability is equivalent to complete $\kappa^{\prime\prime}$-normality of the dual space. Moreover, in [12, Theorem 5.8] they provide an intrinsic characterisation of dualisability in terms of the corresponding non-commutative pointed convex set of the selfadjoint operator space, see [16, Theorem 4.5] of Kennedy, the third author and Manor. We show that when $\kappa=1$ then we have an automatic embedding property. Thus, apart from a good duality framework, in completely approximately 1-generated selfadjoint operator spaces one obtains automatically Arveson’s Extension Theorem.
@@
We note that due to the Hahn-Banach Theorem, every positive functional admits a selfadjoint extension with the same norm. Complete approximate 1-generation imposes that such extensions are automatically positive due to a Jordan decomposition argument. With Example 4.21 we stress that $E$ may be positively generated, but not completely 1-generated, even if $\mathrm{C}^{*}(E)$ is unital (we further illustrate the independence of approximate positive generation, unitality of the generated C*-algebra, and embeddings in Example 5.8).
@@
Therefore the extension of the maps $\phi$ and $\varphi$ are –each on its own– the only obstruction for the embedding property in Example 1.1. Likewise this shows why the inclusion $E_{1}\subseteq\mathbb{C}^{2}$ is an embedding. In this case the map $\varphi$ is the restriction of the compression to the (2,2) entry, while $\phi$ is implemented by the permutation of the basis elements $e_{1}\leftrightarrow e_{2}$.
@@
We denote by $\partial{\mathcal{F}}_{i}$ the terminal object in each category ${\mathcal{F}}_{i}$, $i=1,2,3$ if it exists. Although $\partial{\mathcal{F}}_{3}$ always exists by [7, Theorem 2.25] it is not true that $\partial{\mathcal{F}}_{1}$ or $\partial{\mathcal{F}}_{2}$ exist, or that they are the same if they exist. The example at the beginning of the introduction gives a case where $\partial{\mathcal{F}}_{1}$ and $\partial{\mathcal{F}}_{2}$ do not exist. Blecher–Kirkpatrick–Neal–Werner show in [2, Theorem 2.3] that $\partial{\mathcal{F}}_{1}$ does exist for $E$ being an approximately positively generated selfadjoint operator space. Under the automatic embedding property, in Corollary 6.3 we show that $\partial{\mathcal{F}}_{2}=\partial{\mathcal{F}}_{3}$ when $E$ is completely approximately 1-generated. Moreover, we obtain the following result regarding hyperrigidity.
Theorem E. (Proposition 6.5, Corollary 6.6) Let $E\subseteq\mathrm{C}^{*}(E)$ be a hyperrigid selfadjoint operator space. Then $\mathrm{C}^{*}_{\textup{env}}(E)\cong\mathrm{C}^{*}(E)$ by a canonical *-isomorphism in the following cases:
@@
    $E$ contains a matrix norm-defining approximate order unit $(e_{\lambda})_{\lambda}$ for $\mathrm{C}^{*}(E)$.
@@ 2. Preliminaries
We present some elements of the theory of selfadjoint operator spaces in the sense of [7, 8, 9, 16, 23, 26, 27]. A matrix ordered operator space $E$ is a norm-closed operator space with a completely isometric involution and a family of norm-closed cones $C_{n}\subseteq M_{n}(E)_{sa}$ such that:
@@
A map $\phi\colon E\to F$ is a complete order embedding if it is completely positive and invertible onto its image with a completely positive inverse.
@@
However this does not hold when surjectivity is dropped. Following [16, Section 2], a map $\vartheta\colon E\to F$ is called an embedding if it is a completely isometric complete order embedding, and the unitisation map $\vartheta^{\#}$ is a unital complete order embedding of operator systems; equivalently $\vartheta^{\#}$ is completely isometric. In [26, Corollary 4.9] Werner shows that the unitisation construction recovers the known unitisation when restricted to C*-algebras. We will give an alternative proof to this fact in Corollary 3.6 through our characterisation of embeddings in Theorem 3.3.
@@ 3. Embeddings
In this section we give a characterisation for a completely isometric complete order embedding between selfadjoint operator spaces to be an embedding. Our first goal is to provide a formula that relates the contractive positive functionals of all levels with the distance to the cones of selfadjoint operator spaces. We will need the following proposition which can be deduced from [10, Corollary 4.14]; we give a direct proof that fits in our context.
@@
Theorem 3.3 allows to detect inclusions that are not embeddings. Let $E$ be a selfadjoint operator space inside a C*-algebra $\mathcal{C}$}. If there is an $x\in M_{n}(E)_{sa}$ and a $\varphi\in{\rm CCP}(M_{n}(E),\mathbb{C})$ with $\|\varphi\|=1$ such that $\varphi(x)\notin\overline{{\rm conv}}\{\sigma_{{\mathcal{C}}}(x)\}$, then the inclusion $E\subseteq{\mathcal{C}}$ is not an embedding. Indeed, if the inclusion $E\subseteq{\mathcal{C}}$ were an embedding, then $\varphi$ would have an extension $\widetilde{\varphi}\in{\rm CCP}(M_{n}({\mathcal{C}}),\mathbb{C})$. As $x$ is selfadjoint it follows that $\varphi(x)=\widetilde{\varphi}(x)\in\overline{{\rm conv}}\{\sigma_{{\mathcal{C}}}(x)\}$ (see for example [21, Lemma 2.10]), which is a contradiction.
@@
    Every map $\phi\in{\rm CCP}(E,M_{k}(\mathbb{C}))$ extends to a map $(\widetilde{\phi}\in{\rm CCP}(F,M_{k}(\mathbb{C}))$ with $\widetilde{\phi}\circ\vartheta=\phi$, for every $k\in\mathbb{N}$.
@@
[…]
$\Phi\colon{\rm C}(\mathbb{T})\to{\mathcal{B}}(L);\Phi(\widehat{z}):=U,$
@@ 4. Positive generation
$\overline{E_{+}-E_{+}}=E_{sa}.$
@@
Let $E\subseteq\mathrm{C}^{*}(E)$ be an approximately positively generated selfadjoint operator space such that $\mathrm{C}^{*}(E)$ is unital. Then every completely bounded completely positive map $\phi\colon E\to{\mathcal{B}}(K)$ extends to a (completely bounded) completely positive map $\widetilde{\phi}\colon\mathrm{C}^{*}(E)\to{\mathcal{B}}(K)$.
@@
Remark 4.6.
Note that Theorem 4.5 remains valid if the assumption that $E$ is approximately positively generated is replaced by the weaker assumption that $E$ is separating for $\mathrm{C}^{*}(E)$.
We give two examples of non-separating selfadjoint operator spaces, that generate a unital C*-algebra, but attain positive functionals that have not positive extensions. In Example 5.8 we show that the conditions of approximate positive generation of $E$, unitality of $\mathrm{C}^{*}(E)$, and the inclusion $E\subseteq\mathrm{C}^{*}(E)$ being an embedding are independent.
@@
where $m$ is the normalised Lebesgue measure and $\widehat{f}(0)$ is the zero-th Fourier coefficient. The coordinate function $\widehat{z}$ is in $F$, and thus we have $\mathrm{C}^{*}(F)={\rm C}(\mathbb{T})$. Moreover, $F$ is not separating for ${\rm C}(\mathbb{T})$ since the unital positive map
@@
This is a bounded functional with $\varphi(A)=0$ and $\varphi(B)=-1$. Hence it is (completely) positive on $E$. We claim that $\varphi$ has no (completely) positive extension on $\mathrm{C}^{*}(E)=M_{2}(\mathbb{C})$.
@@
In Example 4.21 we further show that approximate positive generation of $E$ on its own still does not guarantee that the inclusion in $\mathrm{C}^{*}(E)$ is an embedding. The problem is that positive extensions (which exist by Theorem 4.5) may not preserve the norm.
@@
In the third paragraph of the proof of [8, Proposition 2.13], it is used that, in such a context, the inclusion of $E$ inside $\mathrm{C}^{*}(E)$ extends to a unital complete order embedding $E^{\#}\subseteq\mathrm{C}^{*}(E)^{\#}$, thus allowing the extension of a contractive positive functional of $E^{\#}$ to a contractive positive functional of $\mathrm{C}^{*}(E)^{\#}$, and hence from $E$ to $\mathrm{C}^{*}(E)$. In order to show that $E\subseteq\mathrm{C}^{*}(E)$ is an embedding we can show that [8, Lemma 2.11] is an “if and only if” statement.
@@
We now obtain the following proposition.
@@
We say that a selfadjoint operator space $E$ has the automatic embedding property if every completely isometric complete order embedding $\vartheta\colon E\to F$ with $F$ a selfadjoint operator space is automatically an embedding.
In particular we need just to check the completely isometric complete order Hilbertian embeddings.
@@
Thus, we may pick an element $x\in(M_{n}(E)_{sa})_{1}$ such that $\varphi(x)>\|\varphi\|-\varepsilon$, and $x=x_{1}-x_{2}$ with $x_{1},x_{2}\in M_{n}(E)_{+}$ satisfying $\|x_{1}\|,\|x_{2}\|\leq\|x\|$. We then have
@@
Since $\varepsilon$ was arbitrary and $\|x_{1}\|\leq\|x\|\leq 1$ we obtain
@@
We finish the section with an example to demonstrate that the assumption of completely (approximately) 1-generation in Theorem 4.17 cannot be relaxed to mere (approximately) positively generation if one wishes to obtain an automatic embedding property.
@@
Note that $\varphi$ does extend to a positive functional on all of $M_{2}(\mathbb{C})$ by Theorem 4.5, but not with preservation of norm. One such extension is the functional $\widetilde{\varphi}\colon M_{2}(\mathbb{C})\to\mathbb{C}$ given by

### v2

@@ 1.2. The embedding problem
This example exhibits that embeddings should reflect both the extrinsic and the intrinsic aspects of positivity inherited from the C*-cover, e.g., extensions of maps, spectrum, positive decomposition and distance from cones.
@@
Theorem A. ([19], [23], [9], Theorem 3.3, Theorem 3.8) Let $\vartheta\colon E\to F$ be a completely isometric complete order embedding between selfadjoint operator spaces. Then the following are equivalent:
@@
Equivalence [(i) $\Leftrightarrow$ (ii)] is shown in [9, Corollary 2.3]. Equivalence [(i) $\Leftrightarrow$ (iii)] is shown in a broader context in [19, Lemma 4.2 (a)]. Equivalence [(iv) $\Leftrightarrow$ (v)] is shown in [23, Corollary 4.2]. Equivalence [(iii) $\Leftrightarrow$ (iv)] follows by a BW-compactness argument and is provided in Theorem 3.8. Furthermore, equivalence [(i) $\Leftrightarrow$ (v)] is directly shown in Theorem 3.3 with a connection of Russell’s gauge maximality with the description of the positive cones of the unitisation by van Dobben de Bruyn [10]. A key step in this approach is the distance formula
@@ 1.3. The automatic embedding property
Theorem B. (Corollary 4.6, Proposition 5.4) Let $E\subseteq\mathrm{C}^{*}(E)$ be a selfadjoint operator space. Then the following hold:
@@
Approximately positively generated spaces $E$ are in abundance and this extra structure demonstrates strong connections with the ambient C*-algebra. In Proposition 4.3 we combine the notion of rigidity at zero of Salomon [24] with the works of Humeniuk–Kennedy–Manor [12] and of the third author [17] to show that $E$ is approximately positively generated if and only if $M_{n}(E)$ is approximately positively generated for all $n\in\mathbb{N}$. This is an extension of [12, Proposition 8.4] to the approximation setting.
@@
Ng proves in [20, Theorem 3.9] that a selfadjoint operator space $E$ is completely $\kappa$-generated for some $\kappa>0$ if and only if it is completely approximately $\kappa^{\prime}$-generated for some $\kappa^{\prime}>0$ if and only if it is dualisable, i.e., $E^{d}$ is a selfadjoint operator space under the canonical dual matrix cone and with a matrix norm equivalent to the cb-norm. Humeniuk–Kennedy–Manor show in [12, Theorem 5.8] that dualisability is equivalent to complete $\kappa^{\prime\prime}$-normality of the dual space. Moreover, in [12, Theorem 5.8] they provide an intrinsic characterisation of dualisability in terms of the corresponding pointed compact nc-convex set of the selfadjoint operator space, see [16, Theorem 4.5] of Kennedy, the third author and Manor. We show that when $E$ is completely approximately 1-generated then it satisfies the automatic embedding property in the sense of Definition 4.12. Thus, apart from a good duality framework, in completely approximately 1-generated selfadjoint operator spaces one obtains Arveson’s Extension Theorem.
@@
We note that due to the Hahn–Banach Theorem, every positive functional admits a selfadjoint extension with the same norm. Complete approximate 1-generation imposes that such extensions are automatically positive due to a Jordan decomposition argument. With Example 4.22 we stress that $E$ may be positively generated, but not completely 1-generated, even if $\mathrm{C}^{*}(E)$ is unital (we further illustrate the independence of approximate positive generation, unitality of the generated C*-algebra, and embeddings in Example 5.8).
@@
Therefore the extension of the maps $\phi$ and $\varphi$ are –each on its own– the only obstruction for the embedding property in Example 1.1. Likewise this shows why the inclusion $E_{1}\subseteq\mathbb{C}^{2}$ is an embedding. In this case the map $\varphi$ is the restriction of the compression to the (2,2) entry, while $\phi$ is implemented by the unitary permutation of the basis elements.
@@
We denote by $\partial{\mathcal{F}}_{i}$ the terminal object in each category ${\mathcal{F}}_{i}$, $i=1,2,3$ if it exists. Although $\partial{\mathcal{F}}_{3}$ always exists by [7, Theorem 2.25] it is not generally true that $\partial{\mathcal{F}}_{1}$ or $\partial{\mathcal{F}}_{2}$ exist, or that they are the same if they exist. The example at the beginning of the introduction gives a case where $\partial{\mathcal{F}}_{1}$ and $\partial{\mathcal{F}}_{2}$ do not exist. Blecher–Kirkpatrick–Neal–Werner show in [2, Theorem 2.3] that $\partial{\mathcal{F}}_{1}$ does exist for $E$ being an approximately positively generated selfadjoint operator space. Under the automatic embedding property, in Corollary 6.3 we show that $\partial{\mathcal{F}}_{2}=\partial{\mathcal{F}}_{3}$ when $E$ is completely approximately 1-generated. Moreover, we obtain the following result regarding hyperrigidity.
Theorem E. (Proposition 6.5, Corollary 6.6) Let $E\subseteq\mathrm{C}^{*}(E)$ be a hyperrigid selfadjoint operator space. Then $\mathrm{C}^{*}_{\textup{env}}(E)\cong\mathrm{C}^{*}(E)$ by a canonical *-isomorphism if at least one of the following conditions is satisfied:
@@
    $E$ contains a matrix norm-defining approximate order unit $(e_{\lambda})_{\lambda}$ for $E$.
@@ 2. Preliminaries
We present some elements of the theory of selfadjoint operator spaces in the sense of [7, 8, 9, 16, 23, 26, 27]; see also [4, 21] for operator systems theory. A matrix ordered operator space $E$ is a norm-closed operator space with a completely isometric involution and a family of norm-closed cones $C_{n}\subseteq M_{n}(E)_{sa}$ such that:
@@
A map $\phi\colon E\to F$ is a complete order embedding if it is completely positive and invertible onto its image with a completely positive inverse. If $\phi$ is also surjective it is called a complete order isomorphism.
@@
However, $\phi^{\#}$ is not always a unital complete order embedding if $\phi$ is assumed to be merely a completely isometric complete order embedding. Following [16, Section 2], a map $\vartheta\colon E\to F$ is called an embedding if it is a completely isometric complete order embedding, and the unitisation map $\vartheta^{\#}$ is a unital complete order embedding of operator systems; equivalently $\vartheta^{\#}$ is completely isometric. In [26, Corollary 4.9] Werner shows that the unitisation construction recovers the known unitisation when restricted to C*-algebras. We will give an alternative proof to this fact in Corollary 3.6 through our characterisation of embeddings in Theorem 3.3.
@@ 3. Embeddings
In this section we give a characterisation for a completely isometric complete order embedding between selfadjoint operator spaces to be an embedding. Our first goal is to provide a formula that relates the contractive positive functionals of all levels to the distance to the cones of selfadjoint operator spaces. We will need the following proposition which can be deduced from [10, Corollary 4.14]; we give a direct proof that fits in our context.
@@
Theorem 3.3 provides a criterion for detecting inclusions that are not embeddings. Let $E$ be a selfadjoint operator space inside a C*-algebra $\mathcal{C}$}. If there is an $x\in M_{n}(E)_{sa}$ and a $\varphi\in{\rm CCP}(M_{n}(E),\mathbb{C})$ with $\|\varphi\|=1$ such that $\varphi(x)\notin\overline{{\rm conv}}\{\sigma_{{\mathcal{C}}}(x)\}$, then the inclusion $E\subseteq{\mathcal{C}}$ is not an embedding. Indeed, if the inclusion $E\subseteq{\mathcal{C}}$ were an embedding, then $\varphi$ would have an extension $\widetilde{\varphi}\in{\rm CCP}(M_{n}({\mathcal{C}}),\mathbb{C})$. As $x$ is selfadjoint it follows that $\varphi(x)=\widetilde{\varphi}(x)\in\overline{{\rm conv}}\{\sigma_{{\mathcal{C}}}(x)\}$ (see for example [21, Lemma 2.10]), which is a contradiction.
@@
Therefore every completely contractive complete order embedding $\phi\colon E\to{\mathcal{B}}(K)$ is uniquely determined by the contraction $T:=\phi(z)$, for the coordinate function $z$ in ${\rm C}(\mathbb{T})$. By Sz.-Nagy’s dilation theorem, let $U\in{\mathcal{B}}(L)$ be a unitary dilation of $T$ for $L\supseteq K$. By the universal property of the full group C*-algebra $\mathrm{C}^{*}(\mathbb{Z})\cong{\rm C}(\mathbb{T})$ we obtain a unital *-homomorphism
$\Phi\colon{\rm C}(\mathbb{T})\to{\mathcal{B}}(L);\Phi(z):=U,$
@@ 4. Positive generation
$E_{sa}=\overline{E_{+}-E_{+}}.$
In [12, Example 8.1], the authors provide an example of a selfadjoint operator space $E$ that is approximately positively generated, but for which $E_{+}-E_{+}$ is not closed; hence $E$ is not positively generated.
@@
Let $E\subseteq\mathrm{C}^{*}(E)$ be a selfadjoint operator space. If $E$ is separating for $\mathrm{C}^{*}(E)$ and $\mathrm{C}^{*}(E)$ is unital, then every completely bounded completely positive map $\phi\colon E\to{\mathcal{B}}(K)$ extends to a (completely bounded) completely positive map $\widetilde{\phi}\colon\mathrm{C}^{*}(E)\to{\mathcal{B}}(K)$.
@@
By Proposition 4.1 and Theorem 4.5 we obtain the following corollary.
Corollary 4.6.
If $E\subseteq\mathrm{C}^{*}(E)$ is an approximately positively generated selfadjoint operator space and $\mathrm{C}^{*}(E)$ is unital, then every completely bounded completely positive map $\phi\colon E\to{\mathcal{B}}(K)$ extends to a (completely bounded) completely positive map $\widetilde{\phi}\colon\mathrm{C}^{*}(E)\to{\mathcal{B}}(K)$.
We give two examples of non-separating selfadjoint operator spaces, that generate a unital C*-algebra, but attain positive functionals that do not admit positive extensions. In Example 5.8 we show that the conditions of approximate positive generation of $E$, unitality of $\mathrm{C}^{*}(E)$, and the inclusion $E\subseteq\mathrm{C}^{*}(E)$ being an embedding are independent.
@@
where $m$ is the normalised Lebesgue measure and $\widehat{f}(0)$ is the zero-th Fourier coefficient. The coordinate function $z$ is in $F$, and thus we have $\mathrm{C}^{*}(F)={\rm C}(\mathbb{T})$. Moreover, $F$ is not separating for ${\rm C}(\mathbb{T})$ since the unital positive map
@@
This is a bounded functional with $\varphi(E_{11})=0$ and $\varphi(E_{12})=-1$. Hence it is (completely) positive on $E$. We claim that $\varphi$ has no (completely) positive extension on $\mathrm{C}^{*}(E)=M_{2}(\mathbb{C})$.
@@
In Example 4.22 we further show that approximate positive generation of $E$ on its own still does not guarantee that the inclusion in $\mathrm{C}^{*}(E)$ is an embedding. The problem is that positive extensions (which exist by Corollary 4.6) may not preserve the norm.
@@
In order to obtain extensions, the authors in [8] consider selfadjoint operator spaces $E\subseteq\mathrm{C}^{*}(E)$ that contain a matrix norm-defining order unit for the entire $\mathrm{C}^{*}(E)$. In the third paragraph of the proof of [8, Proposition 2.13], it is used that, in this case, the inclusion of $E\subseteq\mathrm{C}^{*}(E)$ extends to a unital complete order embedding $E^{\#}\subseteq\mathrm{C}^{*}(E)^{\#}$, thus allowing the extension of a contractive positive functional. In order to show that $E\subseteq\mathrm{C}^{*}(E)$ is an embedding we can show that [8, Lemma 2.11] is an “if and only if” statement.
@@
To complete the discussion in the context of [8, Section 2.2], we now obtain the following proposition. It will be superseded by Corollary 4.18.
@@
We say that a selfadjoint operator space $E$ has the automatic embedding property if every completely isometric complete order embedding $\vartheta\colon E\to F$ with $F$ a selfadjoint operator space is an embedding.
In particular we just need to check the completely isometric complete order Hilbertian embeddings.
@@
Thus, we may pick an element $x\in(M_{n}(E)_{sa})_{1}$ such that $\varphi(x)>\|\varphi\|-\varepsilon$, and $x=x_{1}-x_{2}$ with $x_{1},x_{2}\in M_{n}(E)_{+}$ satisfying $\|x_{1}\|,\|x_{2}\|\leq 1$. We then have
@@
Since $\varepsilon$ was arbitrary and $\|x_{1}\|\leq 1$ we obtain
@@
As a further application, we improve Proposition 4.11 and require that the matrix norm-defining approximate order unit works just for the selfadjoint operator space, as opposed for the entire generated C*-algebra.
Corollary 4.18.
Let $E$ be a selfadjoint operator space. If $E$ contains a matrix norm-defining approximate order unit $(e_{\lambda})_{\lambda}$ for $E$, then $E$ has the automatic embedding property.
Proof..
It follows from Lemma 4.10 and Corollary 4.16 by noting that $e_{\lambda}^{n}$ is contractive and positive for every $\lambda$ and $n\in\mathbb{N}$. ∎
@@
We finish the section with an example to demonstrate that the assumption of complete (approximate) 1-generation in Theorem 4.17 cannot be relaxed to mere (approximate) positive generation if one wishes to obtain an automatic embedding property.
@@
Note that $\varphi$ does extend to a positive functional on all of $M_{2}(\mathbb{C})$ by Corollary 4.6, but not with preservation of norm. One such extension is the functional $\widetilde{\varphi}\colon M_{2}(\mathbb{C})\to\mathbb{C}$ given by